# Littlewood-Offord bounds on the symmetric groups and applications

### v1

@@ 1.3.2. More general conditions on $v_{i}$
Let $d\geq 2$ be a fixed integer, and let $\delta>0$, $b\neq 0$, and $B>0$ be constants. Suppose that the sequence $(w_{1},\ldots,w_{n})$ satisfies Condition 1.14 for some sufficiently large constant $A>0$. Let $I\subset[n]$ be any subset with $|I|\geq\delta n$, and consider a sequence $(v_{1},\ldots,v_{n})$ that is partially specified by
@@
${\mathbb{P}}\Bigl(\bigl|S_{\pi}-L\bigr|\leq\frac{1}{n}\Bigr)=O_{A}\!\left(\frac{1}{n}\right).$ (9)
@@
${\mathbb{P}}\Bigl(\bigl|S_{\pi}-L\bigr|\leq\frac{1}{n}\Bigr)=O\!\left(\frac{1}{n}e^{-\Theta(L^{2})}\right).$ (10)
@@ 1.3.3. Joint distributions
Let $d\geq 2$ be a fixed integer, and let $\delta>0$, $b\neq 0$, $c\neq 0$, and $B>0$ be given. Suppose that the sequence $(w_{1},\ldots,w_{n})$ satisfies Condition 1.14 for some sufficiently large constant $A>0$. Let $I\subset[n]$ be any subset with $|I|\geq\delta n$, and consider the sequences $(v_{1},\ldots,v_{n})$ and $(v_{1}^{\prime},\ldots,v_{n}^{\prime})$ partially specified by
@@
${\mathbb{P}}\Bigl(\bigl|S_{\pi}({\bm{w}},{\bm{v}})-L_{1}\bigr|\leq\frac{1}{n}\;\wedge\;\bigl|S_{\pi}({\bm{w}},{\bm{v}}^{\prime})-L_{2}\bigr|\leq\frac{1}{n}\Bigr)=O\left(\frac{1}{n^{2}}\right).$ (11)
@@
${\mathbb{P}}\Bigl(\bigl|S_{\pi}({\bm{w}},{\bm{v}})-L_{1}\bigr|\leq\frac{1}{n}\;\wedge\;\bigl|S_{\pi}({\bm{w}},{\bm{v}}^{\prime})-L_{2}\bigr|\leq\frac{1}{n}\Bigr)=O\!\left(\frac{1}{n^{2}}e^{-\Theta(L_{1}^{2}+L_{2}^{2})}\right).$ (12)
@@
It may be possible to extend our approach to sequences of polynomials of non-consecutive degrees; however, we do not pursue this direction in the present paper. We note that the weaker bound
$O\!\left(\frac{1}{n}e^{-\Theta(L_{1}^{2}+L_{2}^{2})}\right),$
involving $O(1/n)$ rather than $O(1/n^{2})$, follows immediately from the one-dimensional estimates. We also refer the reader to Theorem 6.6 for an application of the above result comparing the quantities $S_{\pi}({\bm{w}},{\bm{v}})$ and $S_{\pi}({\bm{w}},{\bm{v}}^{\prime})$.
1.3.4. An inverse theorem via discretization
First, one can obtain an inverse theorem in the continuous setting via a simple discretization argument. For simplicity, assume that $\|{\bm{w}}\|_{\infty}=\|{\bm{v}}\|_{\infty}=1$. Fix a parameter $\alpha>0$ and round each $w_{i}$ and $v_{i}$ to the nearest integer multiple of $\alpha$; denote the resulting vectors by $\bm{w}$$^{\prime}$ and $\bm{v}$$^{\prime}$, respectively. This rounding procedure changes the value of $S_{\pi}$ by at most $3n\alpha$. Moreover, the new sum $S_{\pi}({\bm{w}}^{\prime},{\bm{v}}^{\prime})$ takes values in integer multiples of $\alpha^{2}$.
@@ 1.3.4. An inverse theorem via discretization
For any $\alpha\geq n^{-C^{\prime}}$, there exists a proper symmetric GAP $Q$ of rank $r_{0}=O_{C,C^{\prime},\varepsilon}(1)$ and size $O_{\varepsilon}\bigl(m\rho^{-1}n^{-r_{0}/2}\bigr)$ such that for at least $(1-\varepsilon)n^{4}$ quadruples $(i,j,k,l)$, the quantity $(w_{i}-w_{j})(v_{k}-v_{l})$ is within distance at most $\alpha$ of a point in $Q$.
Next, we introduce a notion–closely related to the Least Common Denominator (LCD) concept of Rudelson and Vershynin [47]–that captures small-ball probabilities rather efficiently.
@@
[60, Theorem 3.2] Under the notation of Definition 1.21, assume that $\|{\bm{u}}\|_{2}\geq n^{3/2}$. Then, for any $\delta\;\geq\;1/\mathbf{LCD}_{\gamma,\kappa}({\bm{w}},{\bm{v}}),$ we have
$\sup_{x\in{\mathbb{R}}}{\mathbb{P}}\big(|S_{\pi}-x|\leq\delta\big)\;\leq\;C\Big(\frac{\delta}{\gamma}\,+\,e^{-\kappa^{2}/2n^{3}}\Big),$
for some absolute constant $C$.
We will include a proof of this result in the appendix for the reader’s convenience.
1.4. Applications
In this subsection, we apply the theorems from the previous two sections to derive several new forward bounds in the discrete setting, as well as applications to the critical points of random polynomials.
1.4.1. Forward theorems
Our first corollary improves Theorems 1.7 and 1.8 under the additional assumption that the multiplicity of any value among the $w_{i}$ is not too close to $n$.
@@ 1.4.1. Forward theorems
Let $0<c<1$ be a constant. Suppose that
$\sup_{y}\#\{i:w_{i}=y\}\leq cn$
and that the $v_{i}$ are all distinct. Then
$\sup_{x}{\mathbb{P}}\bigl(S_{\pi}=x\bigr)=O\!\left(\frac{1}{n^{3/2}}\right),$
where the implied constant may depend on $c$.
Next, we address a problem raised in [2] and discussed in Subsection 1.2.1. We show that if both the $w_{i}$ and the $v_{j}$ are distinct, then the optimal decay rate is $n^{-5/2}$, up to a logarithmic factor. This result may be viewed as an analogue of Theorem 1.2 in the product-space setting.
@@
$\sup_{x}{\mathbb{P}}\bigl(S_{\pi}=x\bigr)=O\!\left(\frac{\log n}{n^{5/2}}\right).$
Note that we allow the possibility that $w_{i}=v_{j}$ for some $i,j$. In particular, taking $w_{i}=v_{i}=i$ yields a bound of order $\Theta(n^{-5/2})$. We will derive these results using Theorem 1.10 in Section 3.
@@
${\mathbb{P}}(S_{\pi}=x)=O\!\left(\frac{1}{n}e^{-\Theta(x^{2})}\right).$ (13)
1.4.2. Critical points of random polynomials
@@
Assuming ${\bm{w}}\neq 0$, Söze [50, Theorem 1] proved that the expected number of nonzero real roots $N_{{\mathbb{R}}}^{\ast}(P_{\pi})$ satisfies
@@
The following condition is analogous to Condition 1.14, but is invariant under shifts and rescaling.
Condition 1.26 (Non-degeneracy: $K$-balanced).
@@
Assume that $\bm{w}$} satisfies Condition 1.26. Then for any nonnegative integer $d$,
@@
1.5. Methods and highlights of our results
To summarize, the highlights of our note include:
  • A nearly optimal inverse result characterizing $a_{ij}$ when $\sup_{x}{\mathbb{P}}(S_{\pi}=x)$ is polynomially large. This result is new and quickly yields several interesting corollaries, such as Theorems 1.23 and 1.24. It also leads to a new conclusion in the setting of the classical inverse result (Theorem 1.6): the support of $\xi$ must possess additional structure. In the proof of Theorem 1.24, we also uncover an unexpected use of the product of differences in additive combinatorics.
  • The establishment of very fine small-ball estimates ${\mathbb{P}}(|S_{\pi}-x|<\delta)$ even when $\delta=o(1/\sqrt{n})$–thus breaking the Berry–Esseen barrier–for various special cases where $(a_{ij})=(w_{i}v_{j})_{1\leq i,j\leq n}$. Our result confirms a heuristic from [50] but also extends far beyond the $O(1/n)$ scale, with optimal decay in the location parameter $L$. This is a new aspect compared to most small-ball estimates in the literature, which usually do not account for $L$.
  • A further highlight of our work is the study of joint distributions of sums induced by $\pi$ (Theorem 1.19), as well as a comparison between the random variables in Theorem 6.6.
  • Although our inverse-type Weyl estimates (e.g., Lemmas 2.6 and 2.7 to be mentioned later) are not entirely new, it is noteworthy that such estimates are required in our analysis of the decay of characteristic functions.
  • Anticipation that our results will lead to numerous interesting applications. As a demonstration, we show that the number of critical points of certain random polynomials is of logarithmic order—a new result we believe to be highly nontrivial.
1.6. Notation
Throughout this paper, we regard $n$ as an asymptotic parameter tending to infinity (in particular, we will implicitly assume that $n$ is larger than any fixed constant, as our claims are all trivial for fixed $n$), and we allow all mathematical objects in the paper to implicitly depend on $n$ unless they are explicitly declared to be “fixed” or “constant.”
We write $X=O(Y)$, $X\ll Y$, or $Y\gg X$ to denote the claim that $|X|\leq CY$ for some fixed constant $C$; this constant $C$ is allowed to depend on other fixed quantities. In the case that $X\ll Y$ and $Y\ll X$, we write $X\asymp Y$ or $X=\Theta(Y)$. We also use $o(Y)$ to denote any quantity bounded in magnitude by $c(n)Y$ for some function $c(n)$ that tends to zero as $n\to\infty$. Again, the function $c(\cdot)$ is permitted to depend on fixed quantities. We denote by $\|x\|_{{\mathbb{R}}/{\mathbb{Z}}}$ the distance from a real number $x$ to the nearest integer. For $x\in{\mathbb{R}}$, we set $e(x):=e^{2\pi ix}$. For a positive integer $n$, we write $n]=\{1,2,\ldots,n\}$.
@@ 2.1. Bound for the characteristic function
By establishing upper bounds on the permanent of such matrices, Roos was able to prove the following result in [45, Theorem 1.4]:
@@
Assume that $S_{\pi}=\sum_{k=1}^{n}a_{k\pi(k)}$. Let
$y_{i,j,k,l}=a_{ik}-a_{jk}-a_{il}+a_{jl},\quad\text{for }i\neq j,\,k\neq l.$
Then the characteristic function $\varphi(t)=\mathbb{E}\,e^{itS_{\pi}}$ satisfies:
  • For any permutation $(l(1),\dots,l(n))$ of $[n]$, we have
    $|\varphi(t)|\leq\prod_{k=1}^{d}\Big(\frac{1}{n(n-1)}\sum_{\begin{subarray}{c}i\neq j\end{subarray}}\cos^{2}\Big(\frac{ty_{i,j,l(2k-1),l(2k)}}{2}\Big)\Big)^{1/2}$
  • and
    $|\varphi(t)|\leq\Big(\frac{1}{n^{2}(n-1)^{2}}\sum_{i\neq j,\,k\neq l}\cos^{2}\Big(\frac{ty_{i,j,k,l}}{2}\Big)\Big)^{d/2},$
where $d=\lfloor n/2\rfloor$.
In this paper we will apply the second conclusion. More precisely, in the proof of Theorem 1.11 we will apply the following corollary.
@@ 2.2. Large deviation result
An immediate consequence of the above lemma is that the moment generating function of the normalized variable
$\bar{S}:=\frac{S_{\pi}-\mathbb{E}S_{\pi}}{\sigma({\bm{w}})\|{\bm{v}}\|_{\infty}}$
is bounded (see, for example, [61, Proposition 2.6.1]):
$m_{\bar{S}}(t):=\mathbb{E}\,e^{t\bar{S}}\leq C_{0}^{\prime}e^{C_{0}^{\prime}t^{2}},\quad\text{for all }t\in{\mathbb{R}},$ (16)
for some constant $C_{0}^{\prime}>0$.
@@ 2.3. Diophantine Properties
For any $0<\delta<1$, there exists a constant $C>0$ such that the following holds. Let $I\subset\{-n,\dots,n\}$ with $|I|\geq\delta n$. Then, for $\frac{C}{n}\leq|b|\leq\frac{1}{C}$ and for any $b_{0}\in{\mathbb{R}}$ we have
$\sum_{r\in I}\|br+b_{0}\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}\asymp_{\delta}n.$
@@
$\sum_{r\in I}\|br+b_{0}\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}\asymp_{C}n.$
@@
Lemma 2.4 is a special case of a more general result concerning polynomial sequences, which we present below.
@@
Let $\delta>0$ and $d\in{\mathbb{Z}}^{+}$ be given. There exists a constant $C>0$ such that the following holds. Let $I\subset\{-n,\dots,n\}$ with $|I|\geq\delta n$. Then, for any $\frac{C}{n}\leq|b|\leq\frac{n^{d-1}}{C}$, we have
$\sum_{r\in I}\left\|\frac{br^{d}+b^{\prime}r^{d-1}+\dots}{n^{d-1}}\right\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}\asymp_{\delta}n,$
@@
To prove this result, we will use a very nice inverse-type Weyl estimate by Tao [54, Corollary 5], which is stated below (and proved in Section B for the reader’s convenience) for positive density form (instead of full density form, $I=\{-n,\ldots,n\}$, as in [54]).
@@
then there is a natural number $q\ll_{d}\delta^{-O_{d}(1)}$ such that
$\|q\alpha_{i}\|_{{\mathbb{R}}/{\mathbb{Z}}}\ll_{d}\delta^{-O_{d}(1)}n^{-i}\quad\text{for all }i=0,\dots,d.$
@@
2.4. Paper Organization
The remainder of the paper is organized as follows. In Section 3, we prove the discrete cases using techniques from [39]. In Section 4, we prove the first continuous result, Theorem 1.15, by first addressing the uniform case with error $O(1/n)$, and then generalizing to the cases $O(1/n^{3/2})$ and $O(1/n^{5/2-o(1)})$; then addressing the $L$-dependent case in the same order, relying on our refined understanding of the characteristic function and its extension to the complex plane. In Section 5, we treat polynomial sequences. In Section 6, we address the most difficult case: proving Theorem 1.19 involving the two-dimensional case. We will do so by relying on techniques developed in both Section 4 and Section 5. Finally, in Section 8, we prove Theorem 1.28. This argument follows a standard sign-change bound (see also [50] and [4]), and leverages Theorem 1.19.
@@ 2.4. Paper Organization
In the forward direction of Theorem 1.11, that $a_{ik}-a_{jk}-a_{il}+a_{jl}\in Q$ for all quadruples, we can easily see that for any $\pi$ and $\pi^{\prime}$, there is a series of $O(n^{2})$ transpositions that move from $S_{\pi}=\sum_{i}a_{i\pi(i)}$ to $S_{\pi^{\prime}}=\sum_{i}a_{i\pi^{\prime}(i)}$, and hence
$\sup_{x}{\mathbb{P}}_{\pi}(S_{\pi}=x)\geq n^{-O(1)}.$
Thus our result claims the converse of this.
Remark 3.1.
In some situations, we can exploit a little further from the conclusions $a_{ik}-a_{jk}-a_{il}+a_{jl}\in Q$ from Theorem 1.11. For instance if we assume that $a_{ik}-a_{jk}-a_{il}+a_{jl}\in Q$, where $Q$ is a symmetric GAP, for all quadruples, then with ${\mathbf{r}}_{i}$ and ${\mathbf{c}}_{j},1\leq i,j\leq n$ being the rows and columns of $(a_{ij})_{1\leq i,j\leq n}$
  1. (1)
    assuming that $k<l$, as $a_{ik}-a_{il}=a_{jk}-a_{jl}$ modulo $Q$ for all $i,j$, we can write c_l = c_k + α_k,l + c_k,l (1,…, 1)^t, where the entries of $\bm{\alpha}_{k,l}$ are in $2Q$, and $c_{k,l}\in{\mathbb{R}}$. As this is true for all $(k,l)$, we can write c_l = c_1 + (_s=1^l-1 c_s, s+1) (1,…, 1)^tmod2nQ.
  2. (2)
    Similarly, r_j = r_i + β_i,j + r_i,j (1,…, 1) where the entries of $\bm{\beta}_{i,j}$ are in $2Q$ and hence r_j = r_1 + (_t=1^j-1 r_t, t+1) (1,…, 1) mod2nQ.
  3. (3)
    Putting together, there are 2(n1)2(n-1) real numbers $r_{t,t+1},c_{s,s+1},1\leq s,t\leq n-1$ such that a_jl = a_11 + _t=1^j-1 r_t,t+1 +_s=1^l-1 c_s,s+1 mod4n Q. In other words, there exist $a_{jl}^{\prime}\in 4nQ$ such that a_jl = a_11 + _t=1^j-1 r_t,t+1 +_s=1^l-1 c_s,s+1 + a_jl’.
@@
For $1\leq i,j\leq n$, let $a_{ij}^{\prime}:=a_{ij}-a_{1j}-a_{i1}+a_{11}\in Q$. Write $S_{\pi}((a_{ij}^{\prime}))=\sum_{i}a_{i\pi(i)}^{\prime}$ for a uniform permutation $\pi$. Then $S_{\pi}((a_{ij}^{\prime}))=S_{\pi}+c$ with $c=na_{11}-\sum_{i=1}^{n}(a_{i1}+a_{1i})$. Since $c$ is independent of $\pi$,
@@
Since $a_{ij}^{\prime}\in Q$, we can write
$a_{ij}^{\prime}=\sum_{s=1}^{r}a_{ij;s}g_{s},\mbox{ for some integers $|a_{ij,s}|\leq N_{s}$}$
and such that
$a_{ik;s}-a_{jk;s}-a_{il;s}+a_{jl;s}\in[-4N_{s},4N_{s}].$ (17)

### v2

@@ 2.3.2. More general conditions on $v_{i}$
Let $d\geq 2$ be a fixed integer, and let $\delta>0$, $b\neq 0$, and $B>0$ be constants. Suppose that the sequence $(w_{1},\ldots,w_{n})$ satisfies Condition 2.5 for some sufficiently large constant $A>0$. Let $I\subset[n]$ be any subset with $|I|\geq\delta n$, and consider a sequence $(v_{1},\ldots,v_{n})$ that is partially specified by
@@
${\mathbb{P}}\Bigl(\bigl|S_{\pi}-L\bigr|\leq\frac{1}{n}\Bigr)=O_{A}\Bigl(\frac{1}{n}\Bigr).$ (9)
@@
${\mathbb{P}}\Bigl(\bigl|S_{\pi}-L\bigr|\leq\frac{1}{n}\Bigr)=O\Bigl(\frac{1}{n}e^{-\Theta(L^{2})}\Bigr).$ (10)
@@ 2.4. Joint distributions
Let $d\geq 2$ be a fixed integer, and let $\delta>0$, $b\neq 0$, $c\neq 0$, and $B>0$ be given. Suppose that the sequence $(w_{1},\ldots,w_{n})$ satisfies Condition 2.5 for some sufficiently large constant $A>0$. Let $I\subset[n]$ be any subset with $|I|\geq\delta n$, and consider the sequences $(v_{1},\ldots,v_{n})$ and $(v_{1}^{\prime},\ldots,v_{n}^{\prime})$ partially specified by
@@
${\mathbb{P}}\Bigl(\bigl|S_{\pi}({\bm{w}},{\bm{v}})-L_{1}\bigr|\leq\frac{1}{n}\;\wedge\;\bigl|S_{\pi}({\bm{w}},{\bm{v}}^{\prime})-L_{2}\bigr|\leq\frac{1}{n}\Bigr)=O\Big(\frac{1}{n^{2}}\Big).$ (11)
@@
${\mathbb{P}}\Bigl(\bigl|S_{\pi}({\bm{w}},{\bm{v}})-L_{1}\bigr|\leq\frac{1}{n}\;\wedge\;\bigl|S_{\pi}({\bm{w}},{\bm{v}}^{\prime})-L_{2}\bigr|\leq\frac{1}{n}\Bigr)=O\Big(\frac{1}{n^{2}}e^{-\Theta(L_{1}^{2}+L_{2}^{2})}\Big).$ (12)
@@
It may be possible to extend our approach to sequences of polynomials of non-consecutive degrees; however, we do not pursue this direction in the present paper. We note that the weaker bound $O(\frac{1}{n}e^{-\Theta(L_{1}^{2}+L_{2}^{2})})$, involving $O(1/n)$ rather than $O(1/n^{2})$, follows immediately from the one-dimensional estimates. We also refer the reader to Theorem 8.6 for an application of the above result comparing the quantities $S_{\pi}({\bm{w}},{\bm{v}})$ and $S_{\pi}({\bm{w}},{\bm{v}}^{\prime})$.
2.5. An inverse theorem via discretization
We can obtain an inverse theorem in the continuous setting via a simple discretization argument. For simplicity, assume that $\|{\bm{w}}\|_{\infty}=\|{\bm{v}}\|_{\infty}=1$. Fix a parameter $\alpha>0$ and round each $w_{i}$ and $v_{i}$ to the nearest integer multiple of $\alpha$; denote the resulting vectors by $\bm{w}$$^{\prime}$ and $\bm{v}$$^{\prime}$, respectively. This rounding procedure changes the value of $S_{\pi}$ by at most $3n\alpha$. Moreover, the new sum $S_{\pi}({\bm{w}}^{\prime},{\bm{v}}^{\prime})$ takes values in integer multiples of $\alpha^{2}$.
@@ 2.5. An inverse theorem via discretization
For any $\alpha\geq n^{-C^{\prime}}$ and $n^{{\varepsilon}}\leq n^{\prime}\leq n$, there exists a proper symmetric GAP $Q$ of rank $r_{0}=O_{C,C^{\prime},{\varepsilon}}(1)$ and size $O_{{\varepsilon}}\bigl(m\rho^{-1}(n^{\prime})^{-r_{0}/2}\bigr)$ such that for at least $(1-n^{\prime}/n)n^{4}$ quadruples $(i,j,k,l)$, the quantity $(w_{i}-w_{j})(v_{k}-v_{l})$ lies within distance at most $\alpha$ of a point in $Q$.
To complete this section, we introduce a notion–closely related to the Least Common Denominator (LCD) concept of Rudelson and Vershynin [51]–that captures small-ball probabilities rather efficiently.
@@
[65, Theorem 3.2] Under the notation of Definition 2.12, assume that $\|{\bm{u}}\|_{2}\geq n^{3/2}$. Then, for any $\delta\;\geq\;1/\mathbf{LCD}_{\gamma,\kappa}({\bm{w}},{\bm{v}}),$ we have
$\sup_{x\in{\mathbb{R}}}{\mathbb{P}}\big(|S_{\pi}-x|\leq\delta\big)=O\Big(\frac{\delta}{\gamma}\,+\,e^{-\kappa^{2}/2n^{3}}\Big).$
We will include a proof of this result in Appendix C for the reader’s convenience.
3. Applications
For a vector ${\bm{v}}=(v_{1},\dots,v_{n})$, define $m_{{\bm{v}}}=\max\limits_{a}\#\{i:1\leq i\leq n,\ v_{i}=a\}$, the maximum multiplicity of an entry of $\bm{v}$}.
3.1. Forward theorems
Our first result is an analog of the Erdős–Littlewood–Offord bound, Theorem 1.1, with the assumption that the multiplicity of any value among the $w_{i}$ and $v_{i}$ are not too close to $n$.
@@ 3.1. Forward theorems
Let $0<c<1$ be a constant. Suppose that $m_{{\bm{w}}},m_{{\bm{v}}}\leq cn$. Then
$\sup_{x}{\mathbb{P}}\bigl(S_{\pi}=x\bigr)=O_{c}\Big(\frac{1}{n^{1/2}}\Big).$
The rate $O(n^{-1/2})$ is optimal. Indeed, for ${\bm{w}}={\bm{v}}=(-1,\dots,-1,1,\dots,1)$, the left-hand side equals $\Theta(n^{-1/2})$.
Our next corollary improves Theorems 1.7 and 1.8 under the additional assumption that the multiplicity of any value among the $w_{i}$ is not too close to $n$.
Corollary 3.2.
Let $0<c<1$ be a constant. Suppose that $m_{{\bm{w}}}\leq cn$ and that the $v_{i}$ are all distinct. Then
$\sup_{x}{\mathbb{P}}\bigl(S_{\pi}=x\bigr)=O_{c}\Big(\frac{1}{n^{3/2}}\Big).$
Next, we address a problem raised in [2] and discussed in Subsection 1.2.1. We show that if both the $w_{i}$ and the $v_{j}$ are distinct, then the optimal decay rate is $n^{-5/2}$, up to a logarithmic factor. This result may be viewed as an analog of Theorem 1.2 in the product-space setting.
@@
$\sup_{x}{\mathbb{P}}\bigl(S_{\pi}=x\bigr)=O\Big(\frac{\log n}{n^{5/2}}\Big).$
We allow the possibility that $w_{i}=v_{j}$ for some $i,j$. In particular, taking $w_{i}=v_{i}=i$ yields a bound of order $n^{-5/2}\log n$. An independent and slightly weaker bound of order $n^{-5/2}(\log n)^{2}$ was posted by Hunter et al. [23] two weeks after the initial upload of this manuscript. Under stronger assumptions on $\bm{v}$} and $\bm{w}$}, our method (see Theorem 2.1) can establish arbitrarily strong bounds on the concentration probability ${\mathbb{P}}(S_{\pi}=x)$, whereas their approach does not seem to yield any bound stronger than $n^{-5/2+o(1)}$. We will derive Corollary 3.2 and Corollary 3.3 using Theorem 2.1 in Section 5.
@@
${\mathbb{P}}(S_{\pi}=x)=O\Big(\frac{1}{n}e^{-\Theta(x^{2})}\Big).$ (13)
3.2. Critical points of random polynomials
@@
Assuming ${\bm{w}}\neq 0$, Söze [54, Theorem 1] proved that the expected number of nonzero real roots $N_{{\mathbb{R}}}^{\ast}(P_{\pi})$ satisfies
@@
The following condition is analogous to Condition 2.5, but is invariant under shifts and rescaling.
Condition 3.5 (Non-degeneracy: $K$-balanced).
@@
Assume that $\bm{w}$} satisfies Condition 3.5. Then for any nonnegative integer $d$,
@@
3.3. Singularity of random row-permutation matrices
Let $d\leq n$ be positive integers such that $\min(d,n-d)=\Omega(n)$. Consider the random $\{0,1\}$-matrix $Q_{n,d}$ whose rows are independent uniform random vectors with exactly $d$ ones. Nguyen [40] proved that for every $C>0$, if $n$ is sufficiently large, then
${\mathbb{P}}\bigl(Q_{n,d}\ \text{is singular}\bigr)\leq n^{-C}.$
This polynomial upper bound was later strengthened by Ferber et al. [15], and subsequently by Tran [65], to an exponential bound. Related analogs for the least singular value were studied by Nguyen–Vu [44], Jain [27], Tran [65], and Jain–Sah–Sawhney [28]; see also [1]. We also refer the reader to Section 12 for related discussion. Above all, all existing approaches rely heavily on the $\{0,1\}$ nature of the matrix entries.
By applying Theorem 2.1, we show that the above model admits a substantial generalization.
Theorem 3.9.
Let $0<{\varepsilon}<1$. For every $C>0$ there exists $n_{0}=n_{0}({\varepsilon},C)$ such that the following holds for all $n\geq n_{0}$. Let ${\bm{v}}=(v_{1},\dots,v_{n})\in{\mathbb{R}}^{n}$ satisfy $m_{{\bm{v}}}\leq(1-{\varepsilon})n$ and $\sum_{i=1}^{n}v_{i}\neq 0$. Let $\pi_{1},\dots,\pi_{n}$ be independent uniformly random permutations of $[n]$, and let $Q_{n\times n}$ be the $n\times n$ matrix whose $i$-th row is $\pi_{i}({\bm{v}})$. Then
${\mathbb{P}}(Q_{n\times n}\text{ is singular})\leq n^{-C}.$
Our proof also extends to perturbed matrices $Q_{n\times n}+F$, where $F$ is a deterministic matrix. It is plausible that the above polynomial bound can be strengthened to an exponential bound (or even smaller when the entries of $\bm{v}$} are sufficiently spread out, i.e., $m_{{\bm{v}}}$ is small), or to the setting where $m_{{\bm{v}}}\leq n-n^{\varepsilon}$, but we do not pursue such refinements here. Our main goal is to provide a clean general framework.
@@ 4.1. Bound for the characteristic function
By establishing upper bounds on the permanent of such matrices, Roos [49, Theorem 1.4] was able to prove the following result:
@@
Let $S_{\pi}=\sum_{k=1}^{n}a_{k\pi(k)}$ and define $y_{i,j,k,l}=a_{ik}-a_{jk}-a_{il}+a_{jl}$ for $i\neq j$ and $k\neq l$. Set $d=\lfloor n/2\rfloor$. Then the characteristic function $\varphi(t)=\mathbb{E}e^{itS_{\pi}}$ satisfies
$|\varphi(t)|\leq\Big(\frac{1}{n^{2}(n-1)^{2}}\sum_{i\neq j,\,k\neq l}\cos^{2}\Big(\frac{ty_{i,j,k,l}}{2}\Big)\Big)^{d/2}.$
The proof of Theorem 2.2 will use the following corollary of this result.
@@ 4.2. Large deviation result
An immediate consequence of the above lemma is the following bound on the moment generating function of the normalized variable
$\bar{S}:=\frac{S_{\pi}-\mathbb{E}S_{\pi}}{\sigma({\bm{w}})\|{\bm{v}}\|_{\infty}}.$
There exists a constant $C_{0}^{\prime}>0$ such that (see, for example, [66, Proposition 2.6.1])
$m_{\bar{S}}(t):=\mathbb{E}e^{t\bar{S}}\leq C_{0}^{\prime}e^{C_{0}^{\prime}t^{2}},\qquad t\in{\mathbb{R}}.$ (16)
@@ 4.3. Diophantine Properties
For any $0<\delta<1$, there exists a constant $C>0$ such that the following holds. Let $I\subset\{-n,\dots,n\}$ with $|I|\geq\delta n$. Then, for $\frac{C}{n}\leq|b|\leq\frac{1}{C}$ and any $b_{0}\in{\mathbb{R}}$,
$\sum_{r\in I}\|br+b_{0}\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}=\Theta_{\delta}(n).$
@@
$\sum_{r\in I}\|br+b_{0}\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}=\Theta_{C}(n).$
@@
Lemma 4.4 is a special case of the following more general result on polynomial sequences.
@@
Let $\delta>0$ and $d\in{\mathbb{Z}}^{+}$ be given. There exists a constant $C>0$ such that the following holds. Let $I\subset\{-n,\dots,n\}$ with $|I|\geq\delta n$. Then, for any $\frac{C}{n}\leq|b|\leq\frac{n^{d-1}}{C}$,
$\sum_{r\in I}\Big\|\frac{br^{d}+b^{\prime}r^{d-1}+\dots}{n^{d-1}}\Big\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}=\Theta_{\delta}(n),$
@@
To prove this result, we will use a very nice inverse-type Weyl estimate by Tao [58, Corollary 5], which is stated below (and proved in Appendix B for the reader’s convenience) for positive density form (instead of full density form, $I=\{-n,\ldots,n\}$, as in [58]).
@@
then there is a natural number $q=O_{d}(\delta^{-O_{d}(1)})$ such that
$\|q\alpha_{i}\|_{{\mathbb{R}}/{\mathbb{Z}}}=O_{d}(\delta^{-O_{d}(1)}n^{-i})\quad\text{for all }i=0,\dots,d.$
@@
We begin by showing that the bound in Theorem 2.2 is sharp.
@@
Let $a_{ij}^{\prime}:=a_{ij}-a_{1j}-a_{i1}+a_{11}$. Since $a_{ij}^{\prime}\in Q$, there exist integers $a_{ij,s}$, bounded by $N_{s}$, such that
$a_{ij}^{\prime}=\sum_{s=1}^{r}a_{ij;s}g_{s}.$ (17)
Write $S_{\pi}((a_{ij}^{\prime}))=\sum_{i}a_{i\pi(i)}^{\prime}$ for a uniform permutation $\pi$. Then $S_{\pi}((a_{ij}^{\prime}))=S_{\pi}+c$ with $c=na_{11}-\sum_{i=1}^{n}(a_{i1}+a_{1i})$. As $c$ is independent of $\pi$,